OpenStudy (anonymous):

8x^3=125 Find the real or imaginary solutions by factoring

OpenStudy (raden):

8x^3=125 8x^3 - 125 = 0 (2x)^3 - 5^3 = 0 (2x - 5)((2x)^2 + 2x * 5 + 5^2) = 0 (2x - 5)(4x^2 + 10x + 25) = 0 make each factor equals zero, then solve for x